OpenStudy (anonymous):

simplify: Radical 48 + Radical 27 and explain

OpenStudy (anonymous):

48 is 12 *4 27 is 9*3

OpenStudy (anonymous):

could i add them?

OpenStudy (saifoo.khan):

\[\sqrt{48} = 4\sqrt3\] \[\sqrt{27} = 3 \sqrt3\] \[4\sqrt3 + 3\sqrt3\] \[7 \sqrt3\]

OpenStudy (agreene):

\[\sqrt{48}=4\sqrt{3}\] \[\sqrt{27}=3\sqrt{3}\] ... saifoo beat me to it :(

OpenStudy (saifoo.khan):

:'(

OpenStudy (anonymous):

thanks guys yall the best:D